EVA Report – May 1st

Sol 9 Crew 193 EVA 16 & 17 Summary
Earth-Date 1May2018

EVA 16

EVA Gold Crew Members – Gold000 (CO Patil-Sabale) / Gold 003 (CE Attig) / Gold005 (CGA Shear)

EVA Objectives – Summary
1. Explore Dinosaur Quarry and Lith Canyon to look for emergency EVA sites
2. Repeated GPS recording experiment.

3. Continue to explore the crew biometrics study while on EVA. Raw data here


Results
Found one emergency shelter at (0518800, 4256500), in Lith Canyon.

Issues/Mitigation

  • This EVA was supposed to begin at 0800, but rains delayed it to 0900. No other issues.

ATVs/Rovers Used

  • ATV 300
  • Honda ATV
  • ATV #350.1


Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
  • Adequate hydration – Yes
  • Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

EVA 17

EVA Gold Crew Members – Gold001 (XO Campbell) / Gold 004 (HSO Pandya) /

EVA Objectives – Summary

  1. Explore Zubrin’s Head and White Rock Canyon.
  2. Continue to explore the crew biometrics study while on EVA / Raw data here
  3. Continued search for emergency shelters

Results
Found an emergency shelter at (0518400, 4247000)

Issues/Mitigation

  • Extreme winds and rain forced us to cut this EVA short.


ATVs/Rovers Used

    • Honda ATV
    • ATV #350.3


Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
  • Adequate hydration – Yes
  • Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

Warm Regards!
Astronomer, (Gold 05), Crew #193

Mars Desert Research Station

Operations Report – May 1st

Crew 193 Operations Report 01May2018

SOL: 9

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10.
Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 7:00 pm to 7:55 am.

Solar— SOC 60 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 220 gallons.

Water (static) – 345 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134437.8

Toilet tank emptied: no
Date last emptied: April 29

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 300, 350.1, 350.3

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0.5 Gallons

# Hours the ATVs were Used today: 3

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.
Summary of suits and radios: Still work to do on two piece suits. Repair of suit chargers still in progress. Other than #10, one piece suits nominal.

Summary of Hab operations: Rear airlock door was pulled on while latched (we accidentally locked a crew-member out). While pulling, one screw came loose; it has been replaced by a screw with more length and girth.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment proceeding.

Summary of RAM operations: None today

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Our garbage is ready to be picked up (both burnable and non burnable). There is also cardboard from sawhorse and wagon boxes. We would also like more ATV fuel (ethanol free gasoline).

Sol Summary – May 1st

Crew 193 Sol Summary Report – 1MAY2018

Sol 9

Summary Title: We Sol-emnly Swear We Are Up to No Good, as per the Mars-rauder’s Map

Author: Shawna Pandya, Health and Safety Officer

Mission Status: Nominally nominal

Sol Activity Summary: The day started off later than usual as the EVA-16 team opted to delay their morning excursion until 0900 in the face of threatening storm clouds to the South and East. The morning EVA then progressed as expected, with the PM team keeping up with usual Hab operations, cleaning schedules, GreenHab upkeep and data collection in the interim. Lunch was a Martian feast of curried kidney beans and brown rice (so much better than chianti and fava beans). The EVA-17 team left shortly thereafter to scout the area just west of Zubrin’s head, wisely choosing to turn back 90min ahead of schedule in anticipation of the weather, luckily only meeting with Martian rain for a slight period at the tail end of their sojourn. The crew astronomer managed to capture one solar prominence and a filament. Hab operations otherwise continued as per usual, with one brave crew member opting to try waterless shampoo, and photo-documenting the results (spoiler alert: said crew member still has all her hair).

Look Ahead Plan: EVA 18 will consist of a Green Valley and Mid-Ridge Planitia exploration, whereas EVA 19 will take up an excursion just Southwest of Kissing Camels Ridge. The team continues to function well under the guidance of the dear Commander. Crew development and team-building will continue by way of ASL, team stories, movies and a good ol’ fashioned round of Shanghai, although to be fair, the latter is probably more likely to result in team-breaking than anything else.

Anomalies in work:

1. The crew astronomer was accidentally locked out for a brief spell, but was let in after pleading his case by way of loud noises, and harbors no grudges (we hope).

Weather: Overcast in the AM, rainy and gusty (a Mars-nado, really) in the PM, resulting in a premature return to base on PM EVA.

Crew Physical Status: "So good it hurts," as per one team member.

EVA: EVA-16 proceeded with an exploration of Lith Canyon to the North, and accrued the usual data regarding emergency shelters and crew comfort, safety, health and well-being. Similarly, EVA-17 proceeded with an exploration just West of Zubrin’s Head, and accrued the same data regarding shelters and crew well-being. EVA-17 was halted prematurely due to incrementally inclement weather, which then became overtly inclement.

Reports filed:

GreenHab Report

Astronomy Report

EVA Request

Journalist Report

Science Report

Sol Summary

Photo Report & Photo of the Day

Operations Report

Reports to be filed:

EVA Report

Support Requested:

1. Yogurt. Please send all the yogurt. The crew’s blood levels of Vitamin Y are critically low, and it makes us sad. 🙁

Journalist Report – May 1st

Journalist Report – Earth-Date 1 May 2018/ Mars Sol 9

Author: Doug Campbell

Crew 193 has rounded the halfway point and can see the finish line of their mission on Mars!

The last two days have included four EVAs to different location of the MDRS campus. Unfortunately, inclement weather has forced the shortening of most recent EVAs. Good judgement by each team has seen them stay out of harm’s way in case of flash flooding caused by a hard downpour in the desert.

Despite the shortened length of the EVAs, each team has explored vast distances and taken incredible photography of the varied landscape surrounding their home base. Mountains have been climbed, valleys have been explored, rocks have been kicked and dunes have been slid down.

Inside the hab, the work continues. Many research projects are wrapping up as the mission nears its end. The water-less dishwasher is showing early promise when compared to standard dish washing methods that consume large amounts of water. The green hab is still living up to its name despite the hot weather and the astronomer is having success finding different elements of the sun to photograph.

As a team, Crew 193 is continuing to grow and learn how to work together to achieve their mission goals. The team remains well fed and in good spirits. They have no doubt that this will be both a successful mission and a valuable learning experience!

Science Report – May 1st

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no new sites were noted as useful for emergency evacuation shelters.

2. Sunspot and Solar Flare Monitoring: One small, shallow prominence was photographed on the Sun. No other features were apparent.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: Both EVAs took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water. Using dirt to clean the dishes seems to have some rapid growing bacteria!

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection. No fogging was ever reported. One helmet fogged today on EVA during a brief rainfall. Once rain stopped, it returned to un-fogged almost immediately.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Astronomy Report – May 1st

Astronomy Report

Name: Eric Shear

Crew: 193
Date: May 1, 2018

Sky Conditions: Mostly clear, occasional clouds

Wind Conditions: Low

Observation Start Time: 1330

Observation End Time: 1400

Summary:

One small, shallow prominence was photographed on the Sun. No other features were apparent.

Photos were taken with the Skyris camera and the best images processed in Autostakkert, Registax, Photoshop, and PowerPoint to colorize and bring out contrast. The results are attached.

Objects Viewed: Sun

Problems Encountered:

One filament was spotted, but was nowhere to be found when viewed through the camera. No amount of tuning could bring it back. The prominence discovered today seems to have developed as quickly.

Greenhab Report – May 1st

GreenHab Report

Date: 1May2018

Environmental control: Ambient with door open

Shade cloth: on

Fan: not used today

Average temperatures: 22.1 ℃

Low: 16.8℃

High: 52.7℃

Hours of supplemental light: None

Changes to crops: Yellowing resolving, plants seem to be in slightly better health.

Daily water usage for crops: 7.75 Gallons

Time(s) of watering for crops: Twice: In the morning at 0800 and at 1730.

Morning research observations: None

Changes to research plants: None

Aquaponics: None

Narrative: Yellowed plants are are turning green again after extra watering. Started singing to them today. Selections included Queen (Greatest Hits). If they rise from the dead overnight, we’ll be on to something.

Support/supplies needed: None